(x – 4)(2x + 7) = 0
 and 
(x – 1)(x + 3) = 32
 
It’s fine if you only answer one, thx.

Grade:8

1 Answers

Arun
6 years ago
Dear student
For 1st question
x-4 = 0 and 2x +7 = 0
Hence
X = 4 or x = -7/2
 
Now
2nd question
x² +2x -3 = 32
x² + 2x -35 = 0
on factoring
(x+7)(x-5) = 0
Hence
x = - 7 or x = 5
 
Regards
Arun (askIITians forum expert)